# Scaling-laws for Large Time-series Models

### v1

@@ A.1 Monash
The Monash dataset has been the default repository of open-source time series data used by the academic community for some time [42]. It contains data from a huge variety of sources and contains a wide variety of characteristics. For this work we exclude series that are either too short or are particularly noisy.888 We found through experimentation that removing very noisy datasets significantly improved training stability. We are then left with a total of 23 different sources which add up to a total of $\sim 500\mathrm{M}$ data points; details are given in Tab. 3.
@@ A.2 Climate
NOAA: We primarily gather data from NOAA high-resolution blended analysis of daily sea surface temperature (SST) which includes both temperature measurements and ice level measurements on a $0.25^{\circ}$ grid worldwide.999 The original data can be found here https://downloads.psl.noaa.gov/Datasets/noaa.oisst.v2.highres/. Weather at different points of the grid are intrinsically correlated, especially on such small grid sizes. We therefore downsample the data by a factor of three by randomly choosing grid points without replacement (we do this independently for each dataset).
@@ A.6 Audio
Appendix B In-Sequence Predictions to Forecasting

### v2

@@ A.1 Monash
The Monash dataset has been the default repository of open-source time series data used by the academic community for some time [47]. It contains data from a huge variety of sources and contains a wide variety of characteristics. For this work we exclude series that are either too short or are particularly noisy.444 We found through experimentation that removing very noisy datasets significantly improved training stability. We are then left with a total of 23 different sources which add up to a total of $\sim 500\mathrm{M}$ data points; details are given in Tab. 2.
@@ A.2 Climate
NOAA: We primarily gather data from NOAA high-resolution blended analysis of daily sea surface temperature (SST) which includes both temperature measurements and ice level measurements on a $0.25^{\circ}$ grid worldwide.555 The original data can be found here https://downloads.psl.noaa.gov/Datasets/noaa.oisst.v2.highres/. Weather at different points of the grid are intrinsically correlated, especially on such small grid sizes. We therefore downsample the data by a factor of three by randomly choosing grid points without replacement (we do this independently for each dataset).
@@ A.6 Audio
A.7 Dataset balancing and pre-processing
Appendix B Training details and compute requirements
We use the AdamW optimizer with a batch size of 512, a cosine learning rate scheduler with a linear warm up of 3000 training steps, and train for a total of $10^{5}$ steps. When training on the entire dataset ($\sim 8$B data points), this equates to roughly two epochs. To reduce computational costs we compute the test loss every $\mathcal{O}(200)$ steps and average over a random 10% of the test data each time.
To produce the results in this paper requires $\mathcal{O}(50-70)$ individual production runs. Apart from the 100M parameter run, these were all carried out on single A100 NVIDIA GPU instances, each taking between 1 and 3 days to complete. As such, overall, the work presented here required $\mathcal{O}(150)$ GPU-days of compute. To host the full dataset, we also required a CPU RAM allocation of approximately 250 GB.
Appendix C Learning-rate and architecture dependence
In Fig. 3 we show the effect of changing the maximum learning rate reached at the end of the warm up. The performance of the model (CRPS) clearly depends on the maximum learning rate, and that dependence is itself a function of parameter count. The dependence on maxmimum learning rate is strong enough that it is possible to get better performance with a smaller model, if the maximum learning rate is too small (or too large) for the larger model. Moreover, for a fixed model size we see a clear optimum learning rate above which models diverge (shown as crosses on Fig. 3). To ensure that we used an optimal maximum learning rate as a function of model size, we fit a power law with a constant offset to the best models (at each parameter size) shown in Fig.  3. In the few cases for the largest models where our power law fit overestimates the optimal maximum learning rate (leading to divergence), we slowly reduce the learning rate until we achieve convergence.
In Figure 4 we show the dependence of model performance on architecture choices, showing that performance is largely invariant to the number of heads, while being only weakly sensitive to the aspect ratio (for aspect ratios $\lesssim 100$).
Appendix D Power-law fits
In 7 we provide the power law fits to teh scaling relations shown in Fig. 1.
Table 7: Power-law fits.
MSE CRPS Log-Likelihood
$\log_{10}(A_{0})$ $B_{0}$ $\log_{10}(A_{0})$ $B_{0}$ $\log_{10}(A_{0})$ $B_{0}$
Number of Parameters, $N_{p}$19.47 0.042 -22.64 0.036 4.33 0.151
Training Compute, $\mathcal{C}$38.88 0.031 -43.03 0.028 -6.65 0.101
Dataset Size, $\mathcal{D}$8.91 0.062 -30.42 0.027 7.00 0.188
Appendix E In-Sequence Predictions to Forecasting